An engine raises 100 kg of water through a height of 60m in 20 seconds. What is the power of the engine(take g as 10ms-2)
1
Expert's answer
2021-08-24T14:42:30-0400

In order to raise "m = 100kg" of water through the height "h = 60m" the following amount of work is required:


"W = mgh"

By definition, the power is:


"P = \dfrac{W}{t} = \dfrac{mgh}{t}"

where "t= 20s" is the time reqired to do the work.

Thus, obtain:


"P = \dfrac{100kg\cdot 10m/s^2\cdot 60m}{20s} = 3000W"

Answer. 3000 W.